A shipment of 15 televisions sets contains 3 defective sets. A hotel purchases 9 of these televisions sets. What is the probability that the hotel receives at least one of the defective sets?
Answer #1

probability that the hotel receives at least one of the defective sets =1-P(no defective sets)

=1-12C9/15C9 =1-220/5005 =1-0.044 =0.9560